800 mmhg how many liters would this same gas occupy at a pressure of 0.75 atm at constant tempeture and omount o gas
1
Expert's answer
2020-05-08T14:11:30-0400

We have T = const and n = const. According to the Boyle's law:


pV = const, from here:

p1V1 = p2V2


For further calculations we need the initial volume of gas (V1). Wr have p1 = 800 mmhg = 1.05 atm and p2 = 0.75 atm.


V2 = p1V1/p2 = V1*1.05/0.75 = 1.4 V1


So, gas with volume V1 at 800 mmHg would occupy volume 1.4 V1 at 0.75 atm.